Sales/total assets = 4x Return on assets = 10% Return on equity = 15%

What is Goldstar's profit margin? What is Goldstar's debt ratio?
Answer #1

Return on Assets = 10%

Net Income/Total assets = 10%

Total Assets = 10*Net Income............(1)

Sales/Total Assets = 4

Total Assets = Sales/4...............(2)

Equating (1) and (2)

Sales = 40*Net Income

Profit Margin = Net Income/Net Sales

= Net income/40*Net Income

= 2.5%

Debt ratio = Debt/Total Assets

Return on Equity = 15%

Net Income/Shareholders Equity = 15%

(Total Assets/10)/(Total Assets – Debt) = 15% -------------------- (Using 1)

Total Assets/10 = 0.15Total Assets – 0.15Debt

Total Assets = 1.5Total Assets – 1.5 Debt

0.5 Total Assets = 1.5Debt

Debt/Total Assets = 0.5/1.5

Debt Ratio = 1/3

Or 33.33%